OpenStudy (kayders1997):

Please check just to make sure my notation is right, the limit of 3x^2g(x) as x approaches 2=? Given that the limit of g(x) as x approaches 2=-6

OpenStudy (kayders1997):

So My notation is 3 on the outside as the limit of x^3g(x) as x approaches 2=3(-6)^3

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

You applied one limit law,\[\large\rm \lim_{x\to a} c f(x)=c\lim_{x\to a} f(x)\]Bringing the constant outside, yes. But you should apply your other limit law as well, the one for dealing with a product,\[\large\rm \lim_{x\to a}f(x)g(x)=\lim_{x\to a}f(x)\cdot \lim_{x\to a}g(x)\]

OpenStudy (kayders1997):

Huh?

OpenStudy (kayders1997):

Ohhhhh yeah!!! Because x^3 and g(x) are being multiplied

zepdrix (zepdrix):

\[\large\rm \lim_{x\to2} (3x^3)g(x)\quad=\quad \lim_{x\to2} (3x^3)\cdot \lim_{x\to2} g(x)\]Yes good good good :)

zepdrix (zepdrix):

And then pull the 3 out after that.

zepdrix (zepdrix):

Or before, since it's multiplication, the ordering of these rules won't matter.

OpenStudy (kayders1997):

Yeah! Ok!!!

zepdrix (zepdrix):

Woah, is it exclamation day? +_+

OpenStudy (kayders1997):

So than it will be 3(2)^3(-6)

OpenStudy (kayders1997):

Yes haha

zepdrix (zepdrix):

Yes

zepdrix (zepdrix):

Woops, I mean YESSSS!!!!!!!!

zepdrix (zepdrix):

XD

OpenStudy (kayders1997):

-144 and yay!!!!